Need help solving logarithm
log(x^2+4)-log(x+2)=2+log(x-2)

Answer by lwsshak3(11628)

log(x^2+4)-log(x+2)=2+log(x-2)
log(x^2+4)-log(x+2)-log(x-2)=2
log(x^2+4)-(log(x+2)+log(x-2))=2
place under single log
log{[(x^2+4)]/[(x+2)(x-2)]}=2
log{[(x^2+4)]/[(x^2-4]}=2
convert to exponential form: base(10) raised to log of number(2)=number[(x^2+4)]/[(x^2-4]
10^2=(x^2+4)/(x^2-4)=100
100x^2-400=x^2+4
99x^2=404
x^2=404/99≈4.08
x≈2.02